OpenStudy (anonymous):

Find the slope of the line that passes through the points (3, -6) and (6, 12). Please help! I am not good with these!:(

OpenStudy (anonymous):

Slope is y2 -y1 over X2-x1 so \[\frac{ 12 - (-6) }{ 6 - 3 } = \frac{ 18 }{ 3 } = \frac{ 6 }{ 1 }\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

If you have any other questions, post here and ill get back to you.

OpenStudy (anonymous):

Find the slope of the line that passes through the points (8, 7) and (11, 7)??

OpenStudy (anonymous):

@jidery

OpenStudy (anonymous):

Lets try it! Remember that slope is: \[\frac{Y2-Y1 }{ X2-X1 }\] And your equation is (8,7) (11,7) Your equation converts to (X1, Y1) (X2, Y2) Try to input your equation into the forumla and let me know what you get.

OpenStudy (anonymous):

(7-7)/(11-8) Correct?

OpenStudy (anonymous):

@jidery

OpenStudy (anonymous):

Lets try it! Remember that slope is: \[\frac{Y2-Y1 }{ X2-X1 }\] And your equation is (8,7) (11,7) Your equation converts to (X1, Y1) (X2, Y2) Try to input your equation into the forumla and let me know what you get.

OpenStudy (anonymous):

Yeah, so it's (7-7)/(11-8) Correct?

OpenStudy (anonymous):

The slope of that line is 0, because \[\frac{ 0 }{ 3 } = 0\]

OpenStudy (anonymous):

Okay, another one. Find the slope of the line that passes through (-9, 5) and (6, 13). and I already know that (13-5)/(6-(-9) = 8/-3 but does it simplify at all? @jidery

OpenStudy (anonymous):

That is as simplified as that can go, it looks like \[-\frac{ 8 }{ 3 }\] The negative sign goes to the outside of the fraction.